eviant:

Use the table provided to find an exponential model ƒ(x) = abx, where ƒ is the population of Arizona, and the input is the number of years since 2004. Approximate to the thousandth’s place.

eviant:

@Vocaloid

Vocaloid:

uh well you have two data points (0, 5739879) and (1,5939292). and the equation y = ab^x, where a is the starting population in 2004 (given), b is the growth rate, and x is the number of years since 2004 in y = ab^x form that's 5939292 = 5739879(b)^1 so you'd just have to solve for b, the growth rate, first then plug b back into y = 5739879(b)^x, leaving x alone

eviant:

ƒ(x) = 5,739,879 • (1.035)x @Vocaloid

Vocaloid:

good